mikewwe13:

Part A What is the x-coordinate in the solution of the system of equations y = - 2/3x and 4x - 9y = -40? Enter your answer in the box. ___________________ Part B What is the solution to the system of equations? {-5 = x - 3y { 11 = -3x + 7y Select to the empty boxes to create the solution (______, _______) -44, -34, -13, -11, -2, -1, 1, 2, 11, 13, 34, 44

Mehek:

You can find the answers by graphing, elimination, and substitution. Which way would you like to do it?

mikewwe13:

graph

Mehek:

Go on here https://www.desmos.com/calculator

Mehek:

And then put those functions in and tell me where the lines intersect

mikewwe13:

substitution is better

Mehek:

Ok then Substitute y = -2/3x for y in the second equation 4x - 9y = -40 4x - 9(-2/3x) = -40 Try simplifying and solving for x

mikewwe13:

y = 4/9x + 40/9 x = -4

Mehek:

Correct so your answer is x = -4

mikewwe13:

and the second question (2, 1) am i correct ?

mikewwe13:

and btw you're beautiful

Mehek:

No it's not that

mikewwe13:

it's not ?

Mehek:

Nope

Mehek:

https://www.desmos.com/calculator Plot it on here and it will show you the graph with the solution

mikewwe13:

i did and it says x = 2, y = 1

mikewwe13:

(1, 2)

Mehek:

It is giving me (1, 2)

Mehek:

Yes that

mikewwe13:

ok thank you and could i call you cutie ?

Mehek:

I don't think that is necessary lol

Mehek:

No problem

mikewwe13:

ok